Per unit Price (Rupees) Quantity Demanded (Thousand MV) Quantity Supplied (Thousand MV)
20 25 14
24 22 16
26 18 18
28 16 20
Requirements:
Keeping in mind the above table, you are required to:
B. Calculate the price elasticity of demand and price elasticity of supply when per unit price of electricity is rupees 24.
1
Expert's answer
2020-05-25T09:11:13-0400

Price elasticity can be defined as the percentage change in the quantity demanded (Qd) or supplied (Qs) and the percentage change in price.

Mathematically, Price elasticity is defined as showed below:





From the question, rupee 24 is used as our target initial price, the corresponding demand and supplied is shown in the table below:

Per unit Price (Rupees) Quantity Demanded (Thousand MV) Quantity Supplied (Thousand MV)

24 22 16

26 18 18

Therefore, it can be deduced that there is an increased in price from 24 to rupee 26; a decrease in quantity demanded from 22 to 18; and an increased in quantity supplied from 16 to 18.


The Price elasticity of demand = [% change in Q.D.] / [% change in Price]

% change in Price= (26 – 24)/24 * 100 = 8.33%

% change in Q.D = (18 - 22)/22 * 100 = -18.19%

% change in Q.S = (18 - 16)/16 * 100 =  12.5%


The Price elasticity of demand (PED)= % change in Q.D. / % change in Price

                                                           = -18.19/8.33

                                                           = -2.18


The Price elasticity of Supply (PES) = % change in Q.S. / % change in Price

                                                           = 12.5 / 8.33

                                                           =1.5

Hence, The price elasticity of demand can be said to be inelastic because, the PED < 1

The price elasticity of Supply can be said to be elastic because, the PES > 1